# MS-IMAP - A Multi-Scale Graph Embedding Approach for Interpretable Manifold Learning

### v4

@@ 1  Output: MI scores for each feature
            Step 4.1: Compute mutual information between feature $\psi_{\mathbf{x}}(i)$ and the pseudo-labels
4            Step 4.2: $MI\_score=\text{MI}(\psi_{\mathbf{x}}(i),c({\psi_{\mathbf{x}}})$)
5            Step 4.3: Append $MI\_score$ to mi_scores
@@
Figure 2: Clustering results on Cora using a subset of features based on feature importance, comparing MS-IMAP embedding space and the input features.
@@ 6.3 Assessing Feature Importance via Embedding Correspondence
We demonstrate the application of MS-IMAP for feature importance by evaluating clustering performance using $k$-means as a function of the number of features, based on the most significant features identified by two methods: the Laplacian score and Mutual Information-based feature selection. We compared the clustering results (using ARI and AMI) when $k$-means is applied to different subsets of embedding features selected by the Laplacian score and MI-based methods, with results obtained by applying the same feature selection methods directly to the original feature space.
Unlike most manifold learning and graph embedding methods, MS-IMAP produces embeddings where the dimensions can be directly mapped to the original feature space, enabling the assessment of individual feature contributions.
On the Cora dataset, Figure 2 shows that for both feature selection methods, the clustering accuracy obtained using the MS-IMAP embedding is significantly higher than when using the raw features across different subsets of features. This improvement is due to the robustness of the MS-IMAP embedding, while still offering interpretability by linking the embedding dimensions to the original features.

### v5

@@ 3      $\text{mi\_scores}=[]$
      Step 4.1: Compute mutual information between feature $\tilde{\psi}_{\mathbf{x}}(i)$ and the pseudo-labels
4      Step 4.2: $MI\_score=\text{MI}(\tilde{\psi}_{\mathbf{x}}(i),c(\tilde{\psi}_{\mathbf{x}})$)
5      Step 4.3: Append MI_scoreMI\_score to mi_scores
@@
(a) Cora dataset (ARI)
(b) Zillions dataset (ARI)
Figure 2: Clustering results using Adjusted Rand Index (ARI) on Cora and Zillions datasets with a subset of features based on feature importance, comparing the MS-IMAP embedding space and the input features.
(a) Cora dataset (AMI)
(b) Zillions dataset (AMI)
Figure 3: Clustering results using Adjusted Mutual Information (AMI) on Cora and Zillions datasets with a subset of features based on feature importance, comparing the MS-IMAP embedding space and the input features.
The Laplacian Score (LS) and Mutual Information (MI)–based feature importance are two complementary approaches to unsupervised feature selection, with LS focusing on geometric smoothness over a graph and MI targeting feature relevance with respect to clustering structure. Laplacian Score (LS) evaluates how smoothly a feature varies over a graph constructed from the data. It is a fully unsupervised, geometry-driven criterion that does not require clustering or pseudo-labels. A feature receives a low score if it varies slowly and consistently across the graph. This makes LS particularly effective for identifying features that align with smooth manifold structure. However, LS also has several limitations. It is not task-aware: it does not assess whether a feature is relevant for a specific downstream objective such as clustering or classification. Moreover, LS can overrate irrelevant but structured noise—for instance, a smoothly varying but non-informative feature may receive a deceptively low score simply because it is smooth.
In contrast, Mutual Information–based feature importance measures how much information a feature provides about cluster assignments. It quantifies the statistical dependency between each feature and the resulting clusters (e.g., from k-means or MS-IMAP embeddings). A key strength of MI is that it is task-aware: it is self-supervised (or can be seen as supervised with respect to pseudo-labels), capturing which features are actually useful for distinguishing between clusters. MI scores are also directly interpretable—a high MI score indicates that the feature contributes to separating groups—and the method is generally robust to irrelevant noise features, which receive low MI if they do not correlate with clustering structure. However, MI is sensitive to the quality of the clustering used to generate pseudo-labels—poor or unstable clustering can diminish its effectiveness.
5 Experimental Results
We evaluated our approach in both synthetic and real datasets in two application domains. First, we demonstrate the use of MS-IMAP embeddings for feature importance—a novel application that highlights the interpretability of our method. Next, we assess the performance of MS-IMAP in clustering tasks using the full embedding space, comparing it against a range of manifold learning-based embedding techniques.
Datasets We study the performance of MS-IMAP compared to other methods for real datasets. We chose a mix of datasets from varied fields: the Census dataset ([14]) is a financial dataset containing information about individuals extracted from the 1994 US Census; the Zilionis dataset ([40]) is a biological dataset containing single-cell sequencing data from different types of cells; the Animals with Attributes (AWA) ([37]) dataset is an image dataset containing images of animals; The Human Activity Recognition (HAR) ([2]) dataset which consists of sensor data collected from subjects performing six different activities:; and the Cora dataset ([30, 24]) is a graph dataset of a citation network consisting of scientific publications. Note that we only use the features without the natural graph associated with Cora (see more information about each dataset in the Appendix).
@@ 5.1 Assessing Feature Importance via Embedding Correspondence
We demonstrate the application of MS-IMAP for feature importance selection by evaluating clustering performance on subsets of the most significant features. Feature selection is performed with respect to both the input space and the MS-IMAP embedding space, using two alternative unsupervised methods: Laplacian Score and Mutual Information-based selection. We compare clustering results obtained by applying \$k\$-means to subsets of features selected from the embedding, against those selected directly from the original feature space using the same methods. Unlike most manifold learning and graph embedding techniques, MS-IMAP produces embeddings in which each dimension corresponds directly to an input feature. This one-to-one mapping enables direct assessment of individual feature contributions, preserving interpretability.
Evaluation metrics: Clustering performance is measured using Adjusted Rand Index (ARI) and Adjusted Mutual Information (AMI). In all experiments, $k$-means is applied in the embedding space. On the Cora and Zillions datasets, Figures 2(a), 2(b), 3(a), and 3(b) show that, for both feature selection methods and for both evaluation metrics, clustering accuracy is consistently higher when using the MS-IMAP embedding compared to the original features—across varying feature subset sizes. This improvement highlights the robustness of the MS-IMAP embedding, while maintaining interpretability through its feature-wise alignment. Additional experiments on the AWA and HAR datasets, presented in the Appendix, further confirm these gains.
5.2 Clustering Output Embedding Space
We evaluate our method by testing its clustering performance using the full output embedding and comparing it to several representative methods, including UMAP, t-SNE, Diffusion Maps, ISOMAP, PHATE ([26]), PaCMAP ([36]), TriMap ([1]) and HeatGeo ([22]). While most of these methods perform dimensionality reduction, we emphasize that dimensionality reduction is a methodological design choice, not an inherent advantage. Many manifold learning techniques default to low-dimensional embeddings (e.g., 2D or 3D) due to computational constraints such as eigendecomposition, yet this often comes at the cost of reduced expressiveness and, particularly for nonlinear methods, diminished interpretability. However, since most baseline manifold learning methods we compare rely on nonlinear dimensionality reduction, we include clustering results based on the embedding space of the Graph Scattering Transform ([16]) for completeness. The Graph Scattering Transform, which is primarily used in graph classification and semi-supervised learning tasks, can be adapted for unsupervised learning via simple vector-based concatenation. We note that the Graph Scattering Transform offers less interpretability than MS-IMAP, which preserves a one-to-one correspondence between input features and embedding dimensions. Additionally, the increased dimensionality of the embeddings produced by the Graph Scattering Transform can lead to higher computational costs. As shown in Table 1, MS-IMAP consistently delivers stable and competitive performance, ranking best or second-best across all benchmarks. In contrast, state-of-the-art methods often underperform or show instability on at least one dataset. In addition, MS-IMAP provides an interpretable embedding through its one-to-one correspondence with the original features—unlike the other competing graph embedding methods.
Runtime and Computational Complexity The computational complexity of Multi-Scale IMAP is of
$O(NDlog(N))$. Experiments were performed on Virtual Server containers with 32 cores Intel Xeon 8259CL running at 2.50Ghz and 256GB of RAM. Our method is practical for handling datasets with millions of points and dozens to hundreds of features within a few hours, making the method well-suited for real-world applications. For example, on the Cancer QC data-set of 48,969 samples and 306 features took 27 mins. Additional details and running times on all datasets are provided in the Appendix. Note that the increase in running times compared to methods such as UMAP is primarily due to our extension of manifold learning to include interpretability.
Data HAR Census Zilionis AWA Cora
Method ARI AMI ARI AMI ARI AMI ARI AMI ARI AMI
[…]
Table 1: Clustering results comparison using ARI and AMI on the Census, Zilionis, AWA, and Cora datasets. Each entry report the mean clustering accuracy and standard deviation. The best performance is bolded.
[…]
Theorem 3 Let $G=(V,\mathbf{W})$ be a connected graph with $n$ vertices and $f\in PW_{\omega}(G)$ for $\lambda_{1}\leq\omega<\lambda_{\text{max}}$. The SGW operator $\psi$ can be constructed such that the set $S$ is a $\Lambda_{\psi}$-set and the set $U=V\setminus S$ is a uniqueness set for any space $PW_{\omega}(G)$ with $\omega<1/\Lambda_{\psi}$ and $\Lambda_{\psi}<\Lambda$ for any $\varphi\in L_{2}(S)$, where $\Lambda$ is the $\Lambda$-set of the Laplacian operator.
The proof is provided in the Appendix (see Theorem 3 in 8.13). Theorem 3 motivates our use of SGW within the proposed framework. Specifically, it shows that SGW operators provide more compact and effective representations of band limited functions in Paley-Wiener spaces on graphs than the traditional Laplacian operator.